OpenStudy (anonymous):

Write the equation of the line that passes through (–2, 1) and is perpendicular to the line 3x – 2y = 5.

OpenStudy (anonymous):

The line 3x-2y = 5 should be re written into slope intercept form. 3x-2y = 5 -2y = -3x+5 y = (3/2)x5/2 Slope is 3/2